If you are saving money with an interest lower than the inflation, you are losing money, but how will it look if you are saving an extreme amount of money? Let's say 1 billion with a 1 % interest, and 2 % inflation. Then you get 10 million a year, and if you only spend a million a year, would you say you are still worse off? Is the answer as simple as yes, but it doesn't matter because you have enough?
1
Expert's answer
2016-09-23T10:21:03-0400
If you are saving money with an interest lower than the inflation, you are losing money even if you are saving an extreme amount of money. If you are saving 1 billion with a 1 % interest and 2 % inflation, then you get 10 million a year but lose 20 million as a result of inflation, so your real value on the account will be 1 billion*(1 + 0.01 - 0.02) = 1*0.99 = 0.99 billion. And even if you only spend a million a year, then you are still worse off, because the real value of money on account decreased, even if you don't spend anything of your interest paid to you.